OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y – 3 = 8/3(x + 2); (–2, 3)

OpenStudy (anonymous):

multiple choice: y + 3 = –3/8(x – 2) y – 3 = –3/8(x + 2) y – 2 = –3/8(x + 3) y + 3 = 3/8(x – 2)

OpenStudy (anonymous):

y-3=-3/8(x+2) A perpendicular Line is the negative reciprocal. Then any point that is positive turns negative, and any point that is negative turns positive.